OpenStudy (scarlettfarra2000):

Working with Polynomials 1. Multiply and simplify. 2. Simplify and leave answers in STANDARD FORM answers will be below

OpenStudy (scarlettfarra2000):

1) \[3x ^{2}y ^{8}\times7x ^{4}y ^{-4}\]

OpenStudy (anonymous):

So what is the property of exponents when you are multiplying exponential terms with like bases

OpenStudy (scarlettfarra2000):

2) part A \[(9x ^{2}+8x-5)-(7x ^{2}+2x-2)\] part B \[2y(2y+3y ^{2})-y ^{2}(3-y)\]

OpenStudy (anonymous):

Well let's do number one for now What is the rule of exponents when multiplying

OpenStudy (scarlettfarra2000):

I'm not sure @Brill

OpenStudy (danjs):

\[\huge 3x ^{2}y ^{8}\times7x ^{4}y ^{-4}\] that is just a bunch of multiplication of everything together. the numbers can combine, and the like bases can combine, powers of each variable...

OpenStudy (danjs):

recall , x^a * x^b = x^(a + b) you add exponents like that , when two terms are multiplied

OpenStudy (danjs):

\[\huge 21 * x^{2+4} * y^{8-4}\]

OpenStudy (danjs):

21*x^6*y^4

OpenStudy (scarlettfarra2000):

so it'll be \[21^{6}\times y ^{4} \]

OpenStudy (danjs):

write that exponent rule down, for this one \[\large (9x ^{2}+8x-5)-(7x ^{2}+2x-2)\] all those terms are being added together, the only thing to change, is since the second parenthesis is - , you have to distribute that -1 into it, changes all the signs around

OpenStudy (danjs):

\[\large 9x ^{2}+8x-5 - 7x ^{2} - 2x +2\]

OpenStudy (danjs):

then combine like terms, the x^2, the x, and the numbers can each combine

OpenStudy (scarlettfarra2000):

like this\[2x ^{2}+6x-7?\]

OpenStudy (danjs):

yeah you got it, cept -5 + 2 is -3 on the end

OpenStudy (scarlettfarra2000):

Thank you so much for your help

OpenStudy (danjs):

this one uses that same rule for adding powers from before, first distribute and expand out those parenthesis terms \[\huge 2y(2y+3y ^{2})-y ^{2}(3-y)\] \[\huge 4y^2+6y ^{3} -3y ^{2}+y^3\]

OpenStudy (danjs):

then just put together the same powers like before